# Sparse Hanson-Wright Inequalities with Applications

### v3

@@ Abstract
We derive new Hanson-Wright-type inequalities tailored to the quadratic forms of random vectors with sparse independent components. Specifically, we consider cases where the components of the random vector are sparse $\alpha$-subexponential random variables with $\alpha>0$. When $\alpha=\infty$, these inequalities can be seen as quadratic generalizations of the classical Bernstein and Bennett inequalities for sparse bounded random vectors. To establish this quadratic generalization, we also develop new Bersntein-type and Bennett-type inequalities for linear forms of sparse $\alpha$-subexponential random variables that go beyond the bounded case $(\alpha=\infty)$. Our proof relies on a novel combinatorial method for estimating the moments of both random linear forms and quadratic forms.
We present two key applications of these new sparse Hanson-Wright inequalities: (1) A local law and complete eigenvector delocalization for sparse $\alpha$-subexponential Hermitian random matrices, generalizing the result of He et al. (2019) beyond sparse Bernoulli random matrices. To the best of our knowledge, this is the first local law and complete delocalization result for sparse $\alpha$-subexponeitial random matrices down to the near-optimal sparsity $p\geq\frac{\mathrm{polylog}(n)}{n}$ when $\alpha\in(0,2)$ as well as for unbounded sparse sub-gaussian random matrices down to the optimal sparsity $p\gtrsim\frac{\log n}{n}.$ (2) Concentration of the Euclidean norm for the linear transformation of a sparse $\alpha$-subexponential random vector, improving on the results of Götze et al. (2021) for sparse sub-exponential random vectors.
11footnotetext: Department of Mathematics, University of California Irvine, Irvine, CA 92697, USA, yiyunh4@uci.edu.22footnotetext: Department of Mathematics, Hong Kong University of Science and Technology, Clear Water Bay, Kowloon, Hong Kong, kewang@ust.hk. 44footnotetext: Department of Mathematics, University of Southern California, 3620 Vermont Avenue Los Angeles, CA 90089, USA, yizhezhu@usc.edu.
@@ 1 Introduction
Hanson and Wright [33, 63] consider the case where $X_{i}$’s are centered sub-gaussian random variables. A random variable $\xi$ is sub-gaussian if $\mathbb{P}(|X-\mathbb{E}(X)|>t)\leq 2\exp(-t^{2}/K^{2})$ for any $t>0$. Its sub-gaussian norm is defined as
$\|\xi\|_{\Psi_{2}}:=\inf\{t>0:\mathbb{E}(\exp(|X|/t)^{2})\leq 2\}<\infty.$
Rudelson and Vershynin [52] provide a modern proof of the Hanson-Wright inequality, improving the dependence on the matrix operator norm in the tail bound from [33, 63]. Specifically, they show that if $\max_{1\leq i\leq n}\|X_{i}\|_{\Psi_{2}}\leq K$. Then for any $t\geq 0$,
@@
The Hanson-Wright inequality has become a fundamental tool in the study of high-dimensional probability [59], random matrix theory [60], and signal processing [39]. However, The inequality (1.1) is limited to sub-gaussian vectors and it does not capture the variance dependence of $X$. If we consider a random vector with centered Ber$(p)$ entries where $p\to 0$ as $n\to\infty$, the inequality (1.1) holds with $K\asymp|\log p|^{-1/2}$ [50]. Given a Bernoulli random vector $X$, the dependence on $p$ of a linear form $\sum_{i=1}^{n}a_{i}X_{i}$ is well-captured by Bennett’s inequality (or Chernoff bound), which plays a fundamental role in various applications in random graph theory and theoretical computer science [4].
These observations naturally lead to the following question: Is there a generalization of Bennett’s inequality for the quadratic forms of sparse random vectors?
@@
Notations: Let $A=(a_{ij})$ be an $n\times n$ real matrix. The Frobenius norm of $A$ is defined as $\|A\|_{F}=\sqrt{\sum_{i,j}a_{ij}^{2}}$, and its operator norm is denoted by $\|A\|$. The maximum norm is given by $\|A\|_{\max}=\max_{i,j}|a_{ij}|$. We also denote $\|A\|_{1,\infty}=\max_{i}\sum_{j}|a_{ij}|$ and $\|A\|_{2,\infty}=\max_{i}(\sum_{j}|a_{ij}|^{2})^{1/2}$. For a vector $a\in\mathbb{R}^{n}$, $\|a\|$ denotes its Euclidean norm and $\|a\|_{\infty}$ denotes the maximum absolute value of its entries. For a real random variable $X$, its $L_{r}$ ($r\geq 1$) norm is denoted by $\|X\|_{L_{r}}=(\mathbb{E}|X|^{r})^{1/r}$. The $\alpha$-Orlicz norm [^1] This is a quasi-norm for $\alpha\in(0,1).$ of a random variable $X$ is defined by
@@
For two functions $f(n),g(n)>0$, we use the asymptotic notations $g(n)\ll f(n)$ and $g(n)=o(f(n))$ if $f(n)/g(n)\to\infty$ as $n\to\infty$. Additionally, we use $f(n)\lesssim g(n)$ or $g(n)\gtrsim f(n)$ to denote that $f(n)\leq Cg(n)$ for some $C>0$. Let $[\log(x)]_{+}$ denote the function equal to $\log(x)$ if $x>1$ and 00 otherwise. In the latter case, we interpret $\frac{1}{[\log(x)]_{+}}=\infty$. For a set $S$, let $\mathbf{1}(S)$ denote the indicator function of $S$.
@@ 1.1 Bernstein-type inequalities for sparse random vectors
Let $X=(X_{1},X_{2},\ldots,X_{n})^{T}$ be a random vector with independent centered sparse $\alpha$-subexponential components, where each $X_{i}$ satisfies $\mathbb{E}(|X_{i}|^{r})\leq p_{i}(Kr^{\frac{1}{\alpha}})^{r}$ for all $r\geq 1$ and some $\alpha\in(0,\infty]$. Le $A$ be a diagonal-free $n\times n$ matrix. Then for any $t>0$,
@@ 1.2 Bennett-type inequalities for sparse random vectors
It is worth noting that for $\alpha>1$, the tail bounds in both Theorem 1 and Theorem 2 are sub-optimal as they do not depend on $\alpha$. To address this limitation, we develop new tail bounds that improve upon both theorems in the large deviation regime for all $\alpha>1$. Below, we present a simplified version of these improved bounds, which has better tail probability estimates compared to Theorems 1 and 2 in the large deviation regime.
@@
Compared to Theorems 1 and 2, when $\alpha>1$, for sufficiently large $t$, the tail probability gains an additional $\log^{1-1/\alpha}(t)$ factor, analogous to the improvement from Bernstein’s inequality to Bennett’s inequality. Notably, for the sparse bounded random variables ($\alpha=\infty$), (1.6) matches Bennett’s inequality in the large deviation regime (up to constants). A detailed comparison is presented in Section 2. The tail behavior described by Bennett’s inequality in the large deviation regime for bounded random variables is known to be sharp (up to universal constants); see [48, Example 2.4]. To the best of our knowledge, these Bennett-type inequalities are new for both the linear form when $\alpha\in(1,\infty)$ and the quadratic form when $\alpha\in(1,\infty]$ of sparse $\alpha$-subexponential random variables.
Before presenting improved precise tail bounds for the linear and quadratic forms for $\alpha$-subexponential random variables with $\alpha\in(1,\infty]$, we introduce several essential notations.
Throughout the context, we denote
$\beta:=1-\frac{1}{\alpha}\quad\text{for }\alpha\in(1,\infty].$ (1.7)
THEOREM 4 (Tail bound for the linear form: $\alpha>1$).
where we denote
REMARK 1.1.
$\exp\left(-C_{\alpha}\frac{t}{KM}\cdot H_{\alpha}(T)\right).$
Next, we present the improved tail bound for the quadratic forms.
THEOREM 5 (Tail bound for the quadratic form: $\alpha>1$).
where we denote
with
REMARK 1.2.
Due to the implicit nature of the Lambert W function, for practical applications of Theorems 4 and 5, we utilize the bounds established in [21]:
$-1-\sqrt{2u}-u<W_{-1}(-e^{-u-1})<-1-\sqrt{2u}-\frac{2}{3}u$ (1.11)
for $u>0$. One can readily verify that Theorem 3 follows directly from applying these bounds (1.11) to Theorems 4 and 5, observing that $W_{-1}\big{(}-\frac{1}{\beta}T^{-\frac{1}{\beta}}\big{)}=W_{-1}\big{(}-e^{-\log\big{(}\frac{\beta}{e}T^{\frac{1}{\beta}}\big{)}-1}\big{)}$.
The proof of Hanson-Wright inequalities for sub-gaussian random vectors in [52] replies on bounding the exponential generating function with a Gaussian comparison argument. Due to the sparsity parameters, a direct adaptation of the proof in [52] will yield a dependence on the sub-gaussian norm of $X_{i}$, which does not have the optimal dependence on $p_{i}$. Moreover, this technique does not apply to $\alpha$-subexponential random variables. Instead, our proof is based on the moment method, which captures the correct variance term in the sub-gaussian tail part and the subexponential tail part also depends on the sparsity parameters. The moment method approach was used in [34, 54, 55]. Our inequality improves the ones in [34] (corresponding to $\alpha=\infty$) to have Bennett-type inequalities for both linear and quadratic forms given in Theorems 4 and 5. We also generalize and improve the inequalities from [54] to sparse $\alpha$-subexponential random vectors for all the regime $\alpha\in(1,\infty]$. A detailed discussion and comparison with the existing literature is presented in Section 2.
@@ 2 Previous results and comparison
Recent progress in the Hanson-Wright-type inequalities and more generally, the polynomials of independent random variables, can be found in [43, 60, 2, 67, 34, 65, 38, 11, 68, 46, 22, 32, 53, 8, 37, 55, 24, 53, 32, 19, 61] and references therein. For the following discussion in this section, let $A$ denote a diagonal-free, symmetric matrix. We mainly focus on the results of concentration for quadratic and linear forms of $\alpha$-subexponential random vectors and sparse random vectors.
@@
Specifically, an extension from sub-gaussian to $\alpha$-subexponential random vectors was obtained in [32, Corollary 1.4 (2)]: for any $\alpha\in(0,1]$, let $X_{1},\dots,X_{n}$ be independent centered random variables with $\|X_{i}\|_{\Psi_{\alpha}}\leq K$. For any $t\geq 0$,
@@
The extension of (2.2) to any $\alpha\in(0,2]$ is discussed in Sambale’s work [53]. Note that for the results mentioned above, their sub-gaussian tail depends on the $\Psi_{\alpha}$-norm of the $X_{i}$, which does not capture the sharp dependence on the sparsity parameter $p$ when $X_{i}$ is sparse.
@@
Now we turn to the results that have a better dependence on $p$ compared to (2.2) for sparse random vectors. The works [67, 51] have focused on quadratic forms involving sparse sub-gaussian random variables. Zhou [67] obtained a result analogous to (2.2) with $\alpha=2$, incorporating $p_{i}$’s only in the sub-gaussian tail, while [51] derived a comparable bound allowing for correlation. In a follow-up work, Zhou [68] examined quadratic forms of non-identical independent Bernoulli variables, bounding the moment generating function of the off-diagonal term. Augeri [8] studied sparse bounded random variables with uniform sparsity $p$ and $\|A\|\leq 1$, obtaining sharp tail bounds in the large deviation regime for $1\gg p\gg\log n/n$ (Lemma 2.8 from [8]). He, Knowles, and Marcozzi [34] established an upper bound on the $L_{r}$ norms $\|X^{T}AX\|_{L_{r}}$ for $X_{i}$’s centered $\mathrm{Ber}(p)$, which proved crucial for analyzing sparse Erdős-Rényi random graphs in the supercritical regime. Their result has been generalized to sparse sub-gaussian random vectors in [9, Lemma 7.4]. While our Theorems 4 and 5 draw inspiration from their work, they ultimately surpass the bounds established in [34].
The study of the concentration of U-statistics also offers insights that are useful for deriving tail estimates of random quadratic forms. In Giné, Latała, and Zine’s study [31], by applying decoupling results and performing straightforward calculations, one can obtain the following tail estimate as a special case of their Theorem 3.3 in [31]: consider $x=(x_{1},\cdots,x_{n})^{T}$ with independent components $x_{i}$ being centered Bernoulli random variables with parameter $p_{i}$. For any $t>0$,
@@
For bounded random variables, the classical Bennett’s inequality [15] (see [17, Theorem 2.9]) provides the following bound: For $\max_{i}X_{i}\leq K$,
@@
where $\nu^{2}=\sum_{i=1}^{n}a_{i}^{2}\mathbb{E}(X_{i}^{2})$ and $h(u)=(1+u)\log(1+u)-u$ for $u>0$. In particular, when $t\gtrsim\frac{\nu^{2}}{K\|a\|_{\infty}}$, the tail bound on the RHS of (2.6) is approximately $\exp\left(-\frac{t}{CK\|a\|_{\infty}}\log\left(\frac{K\|a\|_{\infty}t}{\nu^{2}}\right)\right).$ Compared to our Theorem 3 where $\alpha=\infty$, since $\nu^{2}\leq K^{2}\sum_{i}a_{i}^{2}p_{i}=K^{2}\lambda^{2}$, replacing $\nu^{2}$ with $K^{2}\lambda^{2}$ in this tail bound yields exactly (1.6) up to constants. As shown in [48, Example 2.4], the large deviation behavior captured by Bennett’s inequality achieves optimal sharpness (up to universal constants) for bounded random variables. Improved and refined versions of Bennett’s inequality for the sum of independent bounded random variables have been developed, incorporating the Lambert W function; see, for instance, [66], [36], and [45].
Although an exact analogue of our Theorem 2 for sparse $\alpha$-subexponential random variables has not been found, a similar conclusion might be achievable through existing moment estimates combined with appropriate conditioning techniques. In this direction, we highlight two relevant results from [40]. In [40, Theorem 3.1], the authors leverage results from [42] to establish
@@
where $L_{\alpha}(a)=\|a\|_{\infty}$ if $\alpha\leq 1$ and $L_{\alpha}(a)=\|a\|_{\frac{\alpha}{\alpha-1}}$ if $\alpha>1$. The bounds above are sharp with respect to their dependence on the vector $a$. However, these results do not capture the dependence on the sparsity parameters $p_{i}$. The tail bounds that scale with variance are presented in [40, Theorem 3.2], utilizing proof techniques from [1]:
@@
The factor $(\log(n+1))^{1/\alpha}$ is shown to be sharp by [1, Section 2.2], assuming only $\|X_{i}\|_{\psi_{\alpha}}<\infty$ and variance-based norm bounds. In contrast, our Theorem 2 eliminates this logarithmic term in the subexponential tail by employing a variance proxy $K^{2}\sum_{i}a_{i}^{2}p_{i}$ in the sub-gaussian tail.

### v4

@@ Abstract
We derive new Hanson-Wright-type inequalities tailored to the quadratic forms of random vectors with sparse independent components. Specifically, we consider cases where the components of the random vector are sparse $\alpha$-subexponential random variables with $\alpha>0$. When $\alpha=\infty$, these inequalities can be seen as quadratic generalizations of the classical Bernstein and Bennett inequalities for sparse bounded random vectors. To establish this quadratic generalization, we also develop new Bernstein-type and Bennett-type inequalities for linear forms of sparse $\alpha$-subexponential random variables that go beyond the bounded case $(\alpha=\infty)$. Our proof relies on a novel combinatorial method for estimating the moments of both random linear forms and quadratic forms.
We present two key applications of these new sparse Hanson-Wright inequalities: (1) A local law and complete eigenvector delocalization for sparse $\alpha$-subexponential Hermitian random matrices, generalizing the result of He et al. (2019) beyond sparse Bernoulli random matrices. To the best of our knowledge, this is the first local law and complete delocalization result for sparse $\alpha$-subexponential random matrices down to the near-optimal sparsity $p\geq\frac{\mathrm{polylog}(n)}{n}$ when $\alpha\in(0,2)$ as well as for unbounded sparse sub-gaussian random matrices down to the optimal sparsity $p\gtrsim\frac{\log n}{n}.$ (2) Concentration of the Euclidean norm for the linear transformation of a sparse $\alpha$-subexponential random vector, improving on the results of Götze et al. (2021) for sparse sub-exponential random vectors.
11footnotetext: Department of Mathematics, University of California San Diego, La Jolla, CA 92093, USA,yih130@ucsd.edu22footnotetext: Department of Mathematics, Hong Kong University of Science and Technology, Clear Water Bay, Kowloon, Hong Kong, kewang@ust.hk. 44footnotetext: Department of Mathematics, University of Southern California, 3620 Vermont Avenue Los Angeles, CA 90089, USA, yizhezhu@usc.edu.
@@ 1 Introduction
Hanson and Wright [36, 69] consider the case where $X_{i}$’s are centered sub-gaussian random variables. A random variable $X$ is sub-gaussian if $\mathbb{P}(|X-\mathbb{E}(X)|>t)\leq 2\exp(-t^{2}/K^{2})$ for any $t>0$. Its sub-gaussian norm is defined as
$\|X\|_{\Psi_{2}}:=\inf\{t>0:\mathbb{E}(\exp(|X|/t)^{2})\leq 2\}<\infty.$
Rudelson and Vershynin [58] provide a modern proof of the Hanson-Wright inequality, improving the dependence on the matrix operator norm in the tail bound from [36, 69]. Specifically, they show that if $\max_{1\leq i\leq n}\|X_{i}\|_{\Psi_{2}}\leq K$, then for any $t\geq 0$,
@@
The Hanson-Wright inequality has become a fundamental tool in the study of high-dimensional probability [65], random matrix theory [66], and signal processing [44]. However, the inequality (1.1) is limited to sub-gaussian vectors, and it does not capture the variance dependence of $X$. If we consider a random vector with centered Ber(p)(p) entries where $p\to 0$ as $n\to\infty$, the inequality (1.1) holds with $K\asymp|\log p|^{-1/2}$ [56]. Given a Bernoulli random vector $X$, the dependence on $p$ of a linear form $\sum_{i=1}^{n}a_{i}X_{i}$ is well-captured by Bennett’s inequality (or Chernoff bound), which plays a fundamental role in various applications in random graph theory and theoretical computer science [5]. These observations naturally lead to the following question:
Is there a generalization of Bennett’s inequality for the quadratic forms of sparse random vectors?
@@
We note that one could potentially derive concentration inequalities by first reducing to the special case where $x_{i}\sim\mathrm{Ber}(p_{i})$ and $\xi_{i}$ is a symmetric $\alpha$-Weibull random variable using standard symmetrization and contraction principles (see [46, Theorem 1] for instance). While this approach might lead to simpler proofs, our current approach works directly with the general definition of sparse $\alpha$-subexponential random variables. This generality makes it easier to apply our results in practice.
@@
Notations: Let $A=(a_{ij})$ be an $n\times n$ real matrix. The Frobenius norm of $A$ is defined as $\|A\|_{F}=\sqrt{\sum_{i,j}a_{ij}^{2}}$, and its operator norm is denoted by $\|A\|$. The maximum norm is given by $\|A\|_{\max}=\max_{i,j}|a_{ij}|$. We also denote $\|A\|_{1,\infty}=\max_{i}\sum_{j}|a_{ij}|$ and $\|A\|_{2,\infty}=\max_{i}(\sum_{j}|a_{ij}|^{2})^{1/2}$. For a vector $\in\mathbb{R}^{n}, a\|a\|$ denotes its Euclidean norm and $\|a\|_{\infty}$ denotes the maximum absolute value of its entries. For a real random variable $X$, its $L_{r}$ ($r\geq 1$) norm is denoted by $\|X\|_{L_{r}}=(\mathbb{E}|X|^{r})^{1/r}$. The $\alpha$-Orlicz norm [^1] This is a quasi-norm for $\alpha\in(0,1).$ of a random variable $X$ is defined by
@@
For two functions $f(n),g(n)>0$, we use the asymptotic notations $g(n)\ll f(n)$ and $g(n)=o(f(n))$ if $f(n)/g(n)\to\infty$ as $n\to\infty$. Additionally, we use $f(n)\lesssim g(n)$ or $g(n)\gtrsim f(n)$ to denote that $f(n)\leq Cg(n)$ for some $C>0$. Let $\log(x)]_{+}$ denote the function equal to $\log(x)$ if $x>1$ and 0 otherwise. In the latter case, we interpret $\frac{1}{[\log(x)]_{+}}=\infty$. For a set $S$, let $\mathbf{1}(S)$ denote the indicator function of $S$.
@@ 1.1 Bernstein-type inequalities for sparse random vectors
Let $X=(X_{1},X_{2},\ldots,X_{n})^{T}$ be a random vector with independent centered sparse $\alpha$-subexponential components, where each $X_{i}$ satisfies $\mathbb{E}(|X_{i}|^{r})\leq p_{i}(Kr^{\frac{1}{\alpha}})^{r}$ for all $r\geq 1$ and some $\alpha\in(0,\infty]$. Let $A$ be a diagonal-free $n\times n$ matrix. Then for any $t>0$,
@@ 1.2 Bennett-type inequalities for sparse random vectors
It is worth noting that for $\alpha>1$, the tail bounds in both Theorem 1 and Theorem 2 are sub-optimal as they do not depend on $\alpha$. To address this limitation, we develop new tail bounds that improve upon both theorems in the large deviation regime for all $\alpha>1$. Below, we present a simplified version of these improved bounds, which have better tail probability estimates compared to Theorems 1 and 2 in the large deviation regime. More precise statements of the improved linear form (Theorem 9) and quadratic form (Theorem 10) are deferred to Section 6.
@@
Compared to Theorems 1 and 2, when $\alpha>1$, for sufficiently large $t$, the tail probability gains an additional $\log^{1-1/\alpha}(t)$ factor, analogous to the improvement from Bernstein’s inequality to Bennett’s inequality. Notably, for the sparse bounded random variables ($\alpha=\infty$), (1.6) matches Bennett’s inequality in the large deviation regime (up to constants). A detailed comparison is presented in Section 2. The tail behavior described by Bennett’s inequality in the large deviation regime for bounded random variables is known to be sharp (up to universal constants); see [54, Example 2.4]. To the best of our knowledge, these Bennett-type inequalities are new for both the linear form when $\alpha\in(1,\infty)$ and the quadratic form when $\alpha\in(1,\infty]$ of sparse $\alpha$-subexponential random variables.
The proof of Hanson-Wright inequalities for sub-gaussian random vectors in [58] relies on bounding the exponential generating function with a Gaussian comparison argument. Due to the sparsity parameters, a direct adaptation of the proof in [58] will yield a dependence on the sub-gaussian norm of $X_{i}$, which does not have the optimal dependence on $p_{i}$. Moreover, this technique does not apply to $\alpha$-subexponential random variables. Instead, our proof is based on the moment method, which captures the correct variance term in the sub-gaussian tail part and the subexponential tail part also depends on the sparsity parameters. The moment method approach was used in [37, 60, 61]. Our inequality improves the ones in [37] (corresponding to $\alpha=\infty$) to have Bennett-type inequalities for both linear and quadratic forms given in Theorems 9 and 10. We also generalize and improve the inequalities from [60] to sparse $\alpha$-subexponential random vectors for all the regime $\alpha\in(1,\infty]$. A detailed discussion and comparison with the existing literature is presented in Section 2.
@@ 2 Previous results and comparison
Recent progress in the Hanson-Wright-type inequalities and more generally, the polynomials of independent random variables, can be found in [50, 66, 3, 74, 37, 72, 42, 12, 75, 52, 23, 35, 59, 10, 41, 61, 25, 20, 67] and references therein. For the following discussion in this section, let $A$ denote a diagonal-free, symmetric matrix. We mainly focus on the results of concentration for quadratic and linear forms of $\alpha$-subexponential random vectors and sparse random vectors.
@@
Specifically, an extension from sub-gaussian to $\alpha$-subexponential random vectors was obtained in [35, Corollary 1.4 (2)]: for any $\alpha\in(0,1]$, let $X_{1},\dots,X_{n}$ be independent centered random variables with $\|X_{i}\|_{\Psi_{\alpha}}\leq K$. For any $t\geq 0$,
@@
The extension of (2.2) to any $\alpha\in(0,2]$ is discussed in Sambale’s work [59]. Note that for the results mentioned above, their sub-gaussian tail depends on the $\Psi_{\alpha}$-norm of the $X_{i}$, which does not capture the sharp dependence on the sparsity parameter $p$ when $X_{i}$ is sparse. For related work on moment estimates, two-sided bounds for the symmetric log-convex case ($0<\alpha<1$) have been derived in [43] for general multilinear forms, while the log-concave case ($\alpha\geq 1$) has been studied under various conditions in [48, 50, 1]. The proofs in [35, 59] proceed by converting a quadratic chaos into a bilinear form in two independent copies, after which one typically invokes general multi-level concentration inequalities expressed in terms of $\Psi_{\alpha}$-norms and matrix norms. Such tools are very robust and are well aligned with the regime $0<\alpha\leq 1$. Indeed, in subsequent work [26] we develop a decoupling-based approach that yields sharper sparse Hanson–Wright inequalities for $\alpha\in(0,1].$
For the present paper, our main improvement for $\alpha>1$ is of Bennett type and includes an additional logarithmic gain. This refinement relies on retaining the explicit $p_{i}$-factors in high-moment bounds and optimizing over the moment order; such a gain is typically not captured when sparsity is first summarized through global $\Psi_{\alpha}$-norm control. It would be interesting to understand whether a decoupling framework can be combined with comparably sharp moment bookkeeping to recover the same Bennett-type improvement for $\alpha>1$. We refer to [35] for a detailed discussion of the related results.
@@
Now we turn to the results that have a better dependence on $p$ compared to (2.2) for sparse random vectors. Zhou [74] initiated the study of sparse Hanson-Wright inequality that captures the dependence on $p_{i}$ for sparse sub-Gaussian random vectors $(\alpha=2)$, incorporating $p_{i}$’s in the sub-gaussian tail of (2.2). Notably, when $p_{i}$ are constant, Zhou’s sparse Hanson-Wright inequality recovers (1.1). Some generalizations of Zhou’s inequality can be found in [57, 70]. In a follow-up work, Zhou [75] examined quadratic forms of non-identical independent Bernoulli variables, bounding the moment-generating function of the off-diagonal term. Augeri [10] studied sparse bounded random variables with uniform sparsity $p$ and $\|A\|\leq 1$, obtaining sharp tail bounds in the large deviation regime for $1\gg p\gg\log n/n$ (Lemma 2.8 from [10]). He, Knowles, and Marcozzi [37] established an upper bound on the $L_{r}$ norms $\|X^{T}AX\|_{L_{r}}$ for $X_{i}$’s centered $\mathrm{Ber}(p)$, which proved crucial for analyzing sparse Erdős-Rényi random graphs in the supercritical regime. Their result has been generalized to sparse sub-gaussian random vectors in [9, Lemma 7.4]. While our Theorems 9 and 10 draw inspiration from their work, they ultimately surpass the bounds established in [37].
The study of the concentration of U-statistics also offers insights that are useful for deriving tail estimates of random quadratic forms. In Giné, Latała, and Zinn’s study [33], by applying decoupling results and performing straightforward calculations, one can obtain the following tail estimate as a special case of their Theorem 3.3 in [33]: consider $x=(x_{1},\cdots,x_{n})^{T}$ with independent components $x_{i}$ being centered Bernoulli random variables with parameter $p_{i}$. For any $t>0$,
@@
For bounded random variables, the classical Bennett’s inequality [16] (see [18, Theorem 2.9]) provides the following bound: For $\max_{i}X_{i}\leq K$,
@@
where $\nu^{2}=\sum_{i=1}^{n}a_{i}^{2}\mathbb{E}(X_{i}^{2})$ and $h(u)=(1+u)\log(1+u)-u$ for $u>0$. In particular, when $t\gtrsim\frac{\nu^{2}}{K\|a\|_{\infty}}$, the tail bound on the RHS of (2.6) is approximately $\exp\left(-\frac{t}{CK\|a\|_{\infty}}\log\left(\frac{K\|a\|_{\infty}t}{\nu^{2}}\right)\right).$ Compared to our Theorem 3 where $\alpha=\infty$, since $\nu^{2}\leq K^{2}\sum_{i}a_{i}^{2}p_{i}=K^{2}\lambda^{2}$, replacing $\nu^{2}$ with $K^{2}\lambda^{2}$ in this tail bound yields exactly (1.6) up to constants. As shown in [54, Example 2.4], the large deviation behavior captured by Bennett’s inequality achieves optimal sharpness (up to universal constants) for bounded random variables. Improved and refined versions of Bennett’s inequality for the sum of independent bounded random variables have been developed, incorporating the Lambert W function; see, for instance, [73], [40], and [51].
For sums of independent symmetric $\alpha$-Weibull random variables, two-sided bounds have been established in the literature: Gluskin and Kwapień [34] derived both moment and tail bounds for the sum of $\alpha\geq 1$, while Hitczenko, Montgomery-Smith, and Oleszkiewicz [38] obtained moment inequalities for $\alpha<1$. These results, combined with the symmetrization and contraction principles, enable one to obtain moment estimates for the $\alpha$-subexponential case. Although an exact analogue of our Theorem 2 for sparse $\alpha$-subexponential random variables has not been found, a similar conclusion might be achievable through existing moment estimates combined with appropriate conditioning techniques. In this direction, we highlight two relevant results from [45]. In [45, Theorem 3.1], the authors leverage results from [49] to establish
@@
where $L_{\alpha}(a)=\|a\|_{\infty}$ if $\alpha\leq 1$ and $L_{\alpha}(a)=\|a\|_{\frac{\alpha}{\alpha-1}}$ if $\alpha>1$. The bounds above are sharp with respect to their dependence on the vector $a$. However, these results do not capture the dependence on the sparsity parameters $p_{i}$. The tail bounds that scale with variance are presented in [45, Theorem 3.2], utilizing proof techniques from [2]:
@@
The factor $(\log(n+1))^{1/\alpha}$ is shown to be sharp by [2, Section 2.2], assuming only $\|X_{i}\|_{\psi_{\alpha}}<\infty$ and variance-based norm bounds. In contrast, our Theorem 2 eliminates this logarithmic term in the subexponential tail by employing a variance proxy $K^{2}\sum_{i}a_{i}^{2}p_{i}$ in the sub-gaussian tail.
Technical novelty and comparison of approaches.
Although our proofs are based on the moment method, the improvement in the final tail bounds is not merely a consequence of applying Markov’s inequality to high moments. The key novelty is a graph-structured moment expansion together with a refined counting argument tailored to sparse inputs. Concretely, when expanding $\mathbb{E}\big|\sum_{i\neq j}a_{ij}X_{i}X_{j}\big|^{r}$, each monomial corresponds to a labeled multigraph whose vertices are the indices that appear, and whose edge-multiplicities encode how often each factor $X_{i}X_{j}$ is used. Our estimates exploit this structure by organizing the contribution according to graph features so that each time a new index is activated one incurs an explicit multiplicative penalty through the sparsity parameters $(p_{i})$. This approach is in the spirit of the moment-method analyses in [37, 60, 61], but we leverage the graph structure more systematically and use a sharper enumeration of the relevant combinatorial configurations than [60, 61]. These refinements are what yield the correct variance proxy $\sum_{i\neq j}a_{ij}^{2}p_{i}p_{j}$ and the sparsity-adaptive parameters (e.g. $\gamma_{1,\infty}$ and $\|A\|_{\max}$) in the large-deviation regime.
In contrast to approaches that summarize the inputs through global Orlicz norms and then apply general multi-level polynomial concentration (as discussed above), our moment bounds retain the explicit $p_{i}$-weights throughout the optimization over the moment order. This is essential for our Bennett-type improvements for $\alpha>1$, including the additional logarithmic gain in Theorems 9 and 10.
We conclude the literature review with two open questions. It is unclear to us if a sparsity-adaptive Hanson-Wright inequality exists for Bernoulli random vectors which is optimal for all sparsity regimes. Taking $p_{i}$ to be a constant independent of $n$ in our inequalities does not recover the Hanson-Wright inequality for sub-gaussian random vectors in [58]. For sparse sub-gaussian random vectors, it would be interesting to unify our Theorem 10 and Zhou’s inequality in [74]. After completing this manuscript, we posted a related preprint [26] that further refines sparse Hanson-Wright inequalities for quadratic forms for $\alpha\in(0,2]$. In particular, for $0<\alpha\leq 1$ it yields a sharpened bound that is optimal in several canonical examples, and it also recovers (up to constants) the best-known non-sparse inequalities when $p_{i}\equiv 1$. These results are complementary to the present paper: [26] focuses on $0<\alpha\leq 1$, whereas here we emphasize Bennett-type improvements for $\alpha>1$ (for both linear and quadratic forms), including an additional logarithmic gain in the large-deviation regime obtained via our graph-based moment method.